# Gauri Pradhan

### v1

@@ 1 Introduction
While it has been shown that the current so-called privacy accounting methods for evaluating the privacy bounds for DP-SGD training can produce tight bounds [30], limited empirical studies of the tightness of the privacy bounds for HPO have been presented. There is an example demonstrating privacy leakage through HPO [33], but this involves specifically releasing the hyperparameters in a non-DP setting. Our question focuses on the additional leakage of a released DP model whose hyperparameters have been optimised using training data. In theory, every access to the sensitive training data can increase privacy loss. We test whether this actually leads to higher MIA vulnerability compared to models where HPO has been performed on a different data set from the same population as the training data. Since the both the HPO data and training data are drawn from the same population, the former being a reasonable substitute of the latter for HPO.
@@ 1.1 Our Contributions
  • We show that score-based MIAs is sensitive to the hyperparameters used for training the shadow models and poor choice of hyperparameters may lead to poor attack performance.
@@ 2.1.2 Measuring MIA accuracy
The binary classifier used by the attacker in MIA to predict whether the sample belongs to the training data or not can be used to measure the strength of the MIA. For the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) for this binary classifier as a measure of the vulnerability. Identifying even a small number of samples with high confidence is considered harmful [7] and thus we focus on the regions of small fpr. In some figures, we display the Receiver Operating Characteristic (ROC) which plots the tradeoff between fprs and tprs.
@@ 2.3.2 High utility models under DP
Training high utility models under DP from scratch is challenging. There is a significant gap between training under DP and without DP even for simple computer vision benchmarks [47], resulting in models trained from scratch under DP being unsuitable for real world deployments. The current state-of-the-art results are obtained through transfer learning [53] under the assumption that a public non-sensitive data set can be utilized for pretraining and only the fine-tuning data set is sensitive and needs to be protected through DP.
Prior work has shown that transfer learning under DP is effective for both vision [8, 24, 46, 45] and language tasks [26, 55]. Parameter-efficient fine-tuning using adapters like LoRA [16] and FiLM [36] are competitive in comparison to fine-tuning all parameters under DP as they yield a similar privacy-utility trade-off at a much smaller computational cost [55, 46].
@@ 3 Attacks: Methods
$\texttt{KNOWN}(\mathcal{A},\eta)$ WB-LiRA
@@ 3.1 Strategies for selecting shadow model training hyperparameters
When the target hyperparameters $\eta_{\mathcal{T}}$ are made available together with the target model, the attacker can use the them for training the shadow models. We refer to such an attack as white-box LiRA or WB-LiRA. To implement WB-LiRA, the inputs to Algorithm 1 will be $\{\eta_{\mathcal{T}}\}_{j=1}^{M}$.
@@
3.1.1 Accuracy-based HPO for shadow models (ACC-LiRA)
$\eta_{i}=\textsc{HPO}(\mathcal{D}_{i}).$ (3)
@@ 3.1.1 Accuracy-based HPO for shadow models (ACC-LiRA)
8:        $\phi_{j}\leftarrow\phi_{j}\cup\{\phi_{KL}(\mathcal{N}_{\mathcal{T}}\mid\mid\mathcal{N}_{\mathcal{S}})\}$
@@ 3.1.2 Distribution-based HPO for shadow models (KL-LiRA)
$\phi_{\text{KL}}(\mathcal{N}_{\mathcal{T}}\mid\mid\mathcal{N}_{\mathcal{S}})=\dfrac{1}{2}\Big{[}\dfrac{(\mu_{\mathcal{S}}-\mu_{\mathcal{T}})^{2}}{\sigma_{\mathcal{S}}^{2}}+\dfrac{\sigma_{\mathcal{T}}^{2}}{\sigma_{\mathcal{S}}^{2}}-\text{ln}\dfrac{\sigma_{\mathcal{T}}^{2}}{\sigma_{\mathcal{S}}^{2}}-1\Big{]},$ (4)
@@ 3.2 Computational cost of different attacks
The cost in terms of inference queries is the same for all variants of LiRA. This cost is equal to the number of samples in the union of the shadow data sets $\cup_{i=1}^{N}\mathcal{D}_{i}$. For KL-LiRA, the inference queries on Line $6$ of Algorithm 2 can be answered by using cached responses to these queries. LiRA attack requires no additional computation to run HPO since the target hyperparameters are known to the attacker. Thus its total number of models needed to train is $M$.
@@ 3.3 Testing MIA against a training algorithm
Fig. 3: Overview of attack performances against non-private models. ROC curves depict the performance of different attacks for individual experiments computed over 10 runs. All the plots are generated using 128 shadow models except for FiLM which uses 64 shadow models. The error bars represent the Clopper–Pearson confidence interval associated with the estimated tpr at fixed fpr. WB-LiRA represents a weaker threat model since the attacker has knowledge of the target hyperparameters. MIA vulnerability measured using KL-LiRA matches the vulnerability measured using WB-LiRA. ACC-LiRA’s performs poorly when compared to WB-LiRA and KL-LiRA.
Fig. 4: Overview of relative attack performances against non-private models. MIA vulnerability of KL-LiRA and ACC-LiRA relative to WB-LiRA computed over all the experiments. We normalize the tpr at fixed fpr for KL-LiRA and ACC-LiRA by the corresponding tpr for WB-LiRA for each experiment and run.
@@ 4.1.3 Pretrained Model Architectures
We fine-tuned both ResNet and Vision Transformer architectures pretrained on ImageNet-21K [39]: 1 BiT-M-R50x1 (R-50) [21] with $23.5$M parameters 2 and Vision Transformer ViT-Base-16 (ViT-B) [12] with $85.8$M parameters
@@ 5 Attacks: Results
In this section, we evaluate the performance of different attacks under various threat models. In our experiments, WB-LiRA illustrates the performance of LiRA under the threat model $\texttt{KNOWN}(\mathcal{A},\eta)$, where the attacker knows the target model’s architecture ($\mathcal{A}_{\mathcal{T}}$) and the associated hyperparameters ($\eta_{\mathcal{T}}$). Henceforth, it will serve as the baseline for evaluating the performance of other attacks proposed in this paper, noting that it is a more powerful attack and uses information not available to the other attacks.
@@ 5.1 Testing Attacks in $\texttt{KNOWN}(\mathcal{A})$ Setting
Figure 3 presents the ROC curves depicting the success rate of our attack at different FPRs for each of the experiments. A summary of these plots is provided in Figure 4 which demonstrates the efficacy of these attacks relative to WB-LiRA over all the experiments at different fprs. Figure 3 shows that KL-LiRA consistently performs on par with WB-LiRA across all settings even though the attacker is provided with less information about the training process. On the other hand, ACC-LiRA performs poorly when compared to KL-LiRA and WB-LiRA, especially at very low fprs. At $\textsc{fpr}=0.1\%$, KL-LiRA’s average tpr is near equal to that of WB-LiRA showing that training the shadow models to emulate the output loss distribution of the target model (as done in KL-LiRA) leads to no loss of efficiency for the attack. On the other hand, the average tpr for ACC-LiRA drops to $25\%$ of the corresponding tpr of WB-LiRA. This shows that prioritizing the utility of shadow models yields a subpar attack.
5.2 Testing Attacks in BLACK-BOX Setting
BLACK-BOX represents the most restrictive threat model. The attacker only has access to the final model with no additional information about the training process. Figure 5 demonstrates that the efficiency of ACC-LiRA and KL-LiRA decreases in the BLACK-BOX threat model. Both attacks suffer due to the mismatch between the target and shadow model’s architectures. The effect is particularly significant at low fprs. Their efficacy at $\textsc{fpr}=0.1\%$ degrades by nearly $93\%$ of their performance when target and shadow models are trained with the same architecture. Carlini et al. [7] also show that LiRA perform best when the shadow models’ are trained with the same architecture as the target model. Changing the shadow model’s architecture is expected to alter the performance of the attacks in the BLACK-BOX setting.
Nevertheless, KL-LiRA retains its advantage over ACC-LiRA as a more successful attack. The average tpr at $\textsc{fpr}=0.1\%$ for KL-LiRA in this setting is $75\%$ higher than ACC-LiRA’s.
@@ 5.2 Testing Attacks in BLACK-BOX Setting
5.3 Comparison with Shadow-Model-Free MIAs
5.4 Testing Attacks on DP models
@@
Fig. 7: The plot above summarizes the performance of the attacks against models trained with $(\varepsilon=8,\delta=10^{-5})$-DP as calculated over all the experiments and runs. DP(UB) represents the theoretical upper bound on the tpr at $\varepsilon=8$. For all the experiments, we run the attacks using $M=128$ shadow models except for FiLM which uses $M=64$ shadow models. Using DP offers an effective defence against different versions of LiRA. A detailed version of this plot is available at Figure A.15 which depicts the ROC plot for each experiment.
Figure 8 depicts the relative success of ACC-LiRA and KL-LiRA compared to WB-LiRA when the models are trained with ($\varepsilon=8,\delta=10^{-5}$)-DP. Against models trained with DP, KL-LiRA was found to be a superior attack compared to ACC-LiRA even though its performance suffers relative to WB-LiRA from the non-DP setting. At $\textsc{fpr}=0.1\%$, average tpr for KL-LiRA is 1.4$\times$ higher than ACC-LiRA’s average tpr for $\varepsilon=8$.
Fig. 8: Overview of relative attacker performances against models trained with $(\varepsilon=8,\delta=10^{-5})$-DP. We normalize the MIA vulnerability (tpr) at fixed fpr for KL-LiRA and ACC-LiRA by the corresponding tpr for WB-LiRA for each experiment and run. It can be seen that KL-LiRA is a better choice for attack than ACC-LiRA under DP.
@@ 5.5 Optimizing KL-LiRA
To attack a single target model using KL-LiRA, the attacker needs to find a set of $C$ candidate hyperparameters, $\{\eta_{j}\}_{j=1}^{C}$. This set should be large enough to allow the attacker to find the hyperparameters that best approximate the behaviour of target model’s hyperparameters. We also need to optimize $N$, the number of shadow models that the attacker will need to train for each of the candidate hyperparameters which are used for hyperparameter selection in Algorithm 2. Assuming that the attacker has access to $M$ shadow data sets, a maximally efficient attack would use all the $M$ data sets for the hyperparameter selection process, that is, $C=N=M$. We refer to this as the Full Grid KL-LiRA. However, as mentioned in Section 3.2, the cost for running KL-LiRA scales with $C$ and $N$.
Here, we tried to estimate how many extra models (apart from the $M$ shadow models that would be trained for the attack) will be needed to find the optimal hyperparameters for KL-LiRA to attack one target model such that the attack’s success rate is comparable to its success rate when using the Full Grid. For these experiments, we use $T=20$ HPO trials. Figure 9 shows that for matching the performance of the Full Grid ($C=N=M$), the attackers needs to have $C\geq 16$ candidate hyperparameters and $N\geq 1$ models trained using each of the candidates.
@@ 6 Empirical Privacy Leakage Due To HPO: Methods
Fig. 10: The plot above compares the average MIA vulnerability ($\overline{\textsc{tpr}}$) when the HPO is done using the training data (TD-HPO) against the vulnerability of models for which the HPO is done using external data set (ED-HPO) for different experiments and privacy levels ($\varepsilon=\infty,8,1$) at fixed fpr. The error bars denote the Clopper–Pearson confidence intervals for the tpr along the x and y-axis. If there is no significant additional leakage observed for TD-HPO, the points are expected to fall on the diagonal representing equal privacy leakage for TD- and ED-HPO. If TD-HPO is a more vulnerable setting, the points are expected to fall on the right of the diagonal. Most of the points align with the diagonal showing no significant difference between MIA vulnerability for the two settings.
Next, we study the impact of HPO under DP on the MIA vulnerability of the final released classification model under the assumption that the hyperparameters are not released. A concurrent work by Xiang et al. [50] audits empirical privacy leakage due to DP-HPO in the less restrictive white-box setting.
@@
$\begin{split}\eta^{e}_{i}=\texttt{HPO}(\mathcal{D}^{e}_{i})\\ \mathcal{M}^{e}_{\mathcal{D}_{i},\eta^{e}_{i}}=\texttt{TRAIN}(\mathcal{A},\mathcal{D}_{i},\eta^{e}_{i}),\\ \text{ where }\mathcal{D}^{e}_{i}\cap\mathcal{D}_{j}=\emptyset\quad\forall i,j.\end{split}$ (11)
@@ 7 Empirical Privacy Leakage Due To HPO: Results
Figure 10 summarizes the average MIA vulnerability for both TD-HPO and ED-HPO for different attacks at varying levels of privacy ($\varepsilon=\infty,8,1$) with $\delta=10^{-5}$. The vulnerability to MIA is very similar between TD and ED at all fpr levels, as observed by most of the points in Figure 10 being on or near the diagonal representing $\overline{\textsc{tpr}}_{\text{TD}}=\overline{\textsc{tpr}}_{\text{ED}}$. For the points farther from the diagonal, the error bars (representing the Clopper–Pearson confidence interval) are large and overlap with the diagonal. In Figure 10, We provide a granular version of these plots for each $\varepsilon$ in Figure A.17 and Figure A.18.
@@
Figure 11 shows the $p$-values of different tests and the corresponding BY-adjusted p-values on a logarithmic scale sorted in ascending order.
We can see that the number of rejections of $H_{0}$ after controlling the FDR drops from three to zero for the paired t-test. This suggests that among our experiments, we do not see any examples of TD-HPO resulting in a statistically significant increase of MIA vulnerability compared to the ED-HPO. We observe exactly the same using permutation test: without FDR-control we reject the null three times and after adjusting we retain the null for all the tests (see Figure A.19). Interestingly, no DP experiment yields significant $p$-values even without the adjustment, as all the most significant cases are with non-DP models.

### v2

@@ 1 Introduction
While it has been shown that the current so-called privacy accounting methods for evaluating the privacy bounds for DP-SGD training can produce tight bounds [12], limited empirical studies of the tightness of the privacy bounds for HPO have been presented. There is an example demonstrating privacy leakage through HPO [9], but this involves specifically releasing the hyperparameters in a non-DP setting. A concurrent work by Xiang et al. [13] audits empirical privacy leakage due to DP-HPO in the less restrictive white-box setting. Our question focuses on the additional leakage of a released DP model whose hyperparameters have been optimised using training data. In theory, every access to the sensitive training data can increase privacy loss. We test whether this actually leads to higher MIA vulnerability compared to models where HPO has been performed on a different data set from the same population as the training data. Since the both the HPO data and training data are drawn from the same population, the former being a reasonable substitute of the latter for HPO.
@@ 1.1 Our Contributions
  • We show that score-based MIAs are sensitive to the hyperparameters used for training the shadow models and poor choice of hyperparameters may lead to poor attack performance.
@@ 2.1.2 Measuring MIA accuracy
The binary classifier used by the attacker in MIA to predict whether the sample belongs to the training data or not can be used to measure the strength of the MIA.
For the rest of the paper, we will use the true positive rate (tpr) at a specific false positive rate (fpr) for this binary classifier as a measure of the vulnerability. Identifying even a small number of samples with high confidence is considered harmful [2] and thus we focus on the regions of small fpr. In some figures, we display the Receiver Operating Characteristic (ROC) which plots the tradeoff between fprs and tprs.
@@ 2.3.2 High utility models under DP
Training high utility models under DP from scratch is challenging. There is a significant gap between training under DP and without DP even for simple computer vision benchmarks [22], resulting in models trained from scratch under DP being unsuitable for real world deployments. The current state-of-the-art results are obtained through transfer learning [23] under the assumption that a public non-sensitive data set can be utilized for pre-training and only the fine-tuning data set is sensitive and needs to be protected through DP. Prior work has shown that transfer learning under DP is effective for both vision [24, 25, 26, 27] and language tasks [28, 29]. Parameter-efficient fine-tuning using adapters like LoRA [30] and FiLM [31] are competitive in comparison to fine-tuning all parameters under DP as they yield a similar privacy-utility trade-off at a much smaller computational cost [29, 26].
@@ 3 Attacks: Methods
$\texttt{KNOWN}(\mathcal{A},\eta)$ LiRA
@@ 3.1 Strategies for Selecting Shadow Model Training Hyperparameters
When the target hyperparameters $\eta_{\mathcal{T}}$ are made available together with the target model, the attacker can use the them for training the shadow models as done by Carlini et al. [2] for LiRA.To implement LiRA, the inputs to Algorithm 1 will be $\{\eta_{\mathcal{T}}\}_{j=1}^{M}$.
3.1.1 Accuracy-based HPO for shadow models (ACC-LiRA)
$\eta_{i}=\textsc{HPO}(\mathcal{D}_{i}).$ (3)
@@ 3.1.1 Accuracy-based HPO for shadow models (ACC-LiRA)
8:        $\phi_{j}\leftarrow\phi_{j}\cup\{\phi_{KL}(\mathcal{N}_{\mathcal{T}}||\mathcal{N}_{\mathcal{S}})\}$
@@ 3.1.2 Distribution-based HPO for shadow models (KL-LiRA)
$\phi_{\text{KL}}(\mathcal{N}_{\mathcal{T}}||\mathcal{N}_{\mathcal{S}})=\dfrac{1}{2}\Big{[}\dfrac{(\mu_{\mathcal{S}}-\mu_{\mathcal{T}})^{2}}{\sigma_{\mathcal{S}}^{2}}+\dfrac{\sigma_{\mathcal{T}}^{2}}{\sigma_{\mathcal{S}}^{2}}-\text{ln}\dfrac{\sigma_{\mathcal{T}}^{2}}{\sigma_{\mathcal{S}}^{2}}-1\Big{]},$ (4)
@@ 3.2 Computational Cost of Different Attacks
The cost in terms of inference queries is the same for all variants of LiRA. This cost is equal to the number of samples in the union of the shadow data sets $\cup_{i=1}^{N}\mathcal{D}_{i}$. For KL-LiRA, the inference queries on Line $6$ of Algorithm 2 can be answered by using cached responses to these queries.
LiRA attack requires no additional computation to run HPO since the target hyperparameters are known to the attacker. Thus its total number of models needed to train is $M$.
@@ 3.3 Testing MIA Against a Training Algorithm
Fig. 3: Overview of attack performances against non-private models. ROC curves depict the performance of different attacks for individual experiments computed over 10 runs. All the plots are generated using 128 shadow models except for FiLM which uses 64 shadow models. The error bars represent the Clopper–Pearson confidence interval associated with the estimated tpr at fixed fpr. LiRA represents a weaker threat model since the attacker has knowledge of the target hyperparameters. MIA vulnerability measured using KL-LiRA matches the vulnerability measured using LiRA. ACC-LiRA’s performs poorly when compared to LiRA and KL-LiRA.
Fig. 4: Overview of relative attack performances against non-private models. MIA vulnerability of KL-LiRA and ACC-LiRA relative to LiRA computed over all the experiments. We normalize the tpr at fixed fpr for KL-LiRA and ACC-LiRA by the corresponding tpr for LiRA for each experiment and run.
@@ 4.1.3 Pretrained Model Architectures
We fine-tuned both ResNet and Vision Transformer architectures pretrained on ImageNet-21K [38]: 1 BiT-M-R50x1 (R-50) [39] with $23.5$M parameters 2 and Vision Transformer ViT-Base-16 (ViT-B) [40] with $85.8$M parameters. We briefly discuss the attacking models with relatively weaker pretrained architectures in Section A.2.
@@ 5 Attacks: Results
In this section, we evaluate the performance of different attacks under various threat models. In our experiments, LiRA illustrates the performance of LiRA under the threat model $\texttt{KNOWN}(\mathcal{A},\eta)$, where the attacker knows the target model’s architecture ($\mathcal{A}_{\mathcal{T}}$) and the associated hyperparameters ($\eta_{\mathcal{T}}$). Henceforth, it will serve as the baseline for evaluating the performance of other attacks proposed in this paper, noting that it is a more powerful attack and uses information not available to the other attacks.
@@ 5.1 Testing Attacks in $\texttt{KNOWN}(\mathcal{A})$ Setting
Figure 3 presents the ROC curves depicting the success rate of our attack at different FPRs for each of the experiments. A summary of these plots is provided in Figure 4 which demonstrates the efficacy of these attacks relative to LiRA over all the experiments at different fprs. Figure 3 shows that KL-LiRA consistently performs on par with LiRA across all settings even though the attacker is provided with less information about the training process. On the other hand, ACC-LiRA performs poorly when compared to KL-LiRA and LiRA, especially at very low fprs. At $\textsc{fpr}=0.1\%$, KL-LiRA’s average tpr is near equal to that of LiRA showing that training the shadow models to emulate the output loss distribution of the target model (as done in KL-LiRA) leads to no loss of efficiency for the attack. On the other hand, the average tpr for ACC-LiRA drops to $25\%$ of the corresponding tpr of LiRA. This shows that prioritizing the utility of shadow models yields a subpar attack.
5.2 Comparison with Shadow-Model-Free MIAs
@@
5.3 Testing Attacks in BLACK-BOX Setting
BLACK-BOX represents the most restrictive threat model. The attacker only has access to the final model with no additional information about the training process. Figure 6 demonstrates that the efficiency of ACC-LiRA and KL-LiRA decreases in the BLACK-BOX threat model. Both attacks suffer due to the mismatch between the target and shadow model’s architectures. The effect is particularly significant at low fprs. Their efficacy at $\textsc{fpr}=0.1\%$ degrades by nearly $93\%$ of their performance when target and shadow models are trained with the same architecture. Carlini et al. [2] also show that LiRA perform best when the shadow models’ are trained with the same architecture as the target model. Changing the shadow model’s architecture is expected to alter the performance of the attacks in the BLACK-BOX setting. Nevertheless, KL-LiRA retains its advantage over ACC-LiRA as a more successful attack. The average tpr at $\textsc{fpr}=0.1\%$ for KL-LiRA in this setting is $75\%$ higher than ACC-LiRA’s.
5.4 Testing Attacks on Differentially Private Models
Fig. 7: The plot above summarizes the performance of the attacks against models trained with $(\varepsilon=8,\delta=10^{-5})$-DP as calculated over all the experiments and runs. DP(UB) represents the theoretical upper bound on the tpr at $\varepsilon=8$. For all the experiments, we run the attacks using $M=128$ shadow models except for FiLM which uses $M=64$ shadow models. Using DP offers an effective defence against different versions of LiRA. A detailed version of this plot is available at Figure A4 which depicts the ROC plot for each experiment.
Fig. 8: Overview of relative attacker performances against models trained with $(\varepsilon=8,\delta=10^{-5})$-DP. We normalize the MIA vulnerability (tpr) at fixed fpr for KL-LiRA and ACC-LiRA by the corresponding tpr for LiRA for each experiment and run. It can be seen that KL-LiRA is a better choice for attack than ACC-LiRA under DP.
@@ 5.4 Testing Attacks on Differentially Private Models
Figure 8 depicts the relative success of ACC-LiRA and KL-LiRA compared to LiRA when the models are trained with ($\varepsilon=8,\delta=10^{-5}$)-DP. Against models trained with DP, KL-LiRA was found to be a superior attack compared to ACC-LiRA even though its performance suffers relative to LiRA from the non-DP setting. At $\textsc{fpr}=0.1\%$, average tpr for KL-LiRA is 1.4$\times$ higher than ACC-LiRA’s average tpr for $\varepsilon=8$.
@@ 5.5 Optimizing KL-LiRA
To attack a single target model using KL-LiRA, the attacker needs to find a set of $C$ candidate hyperparameters, $\{\eta_{j}\}_{j=1}^{C}$. This set should be large enough to allow the attacker to find the hyperparameters that best approximate the behaviour of target model’s hyperparameters. We also need to optimize $N$, the number of shadow models that the attacker will need to train for each of the candidate hyperparameters which are used for hyperparameter selection in Algorithm 2. Assuming that the attacker has access to $M$ shadow data sets, a maximally efficient attack would use all the $M$ data sets for the hyperparameter selection process, that is, $C=N=M$. We refer to this as the Full Grid KL-LiRA. However, as mentioned in Section 3.2, the cost for running KL-LiRA scales with $C$ and $N$. Here, we tried to estimate how many extra models (apart from the $M$ shadow models that would be trained for the attack) will be needed to find the optimal hyperparameters for KL-LiRA to attack one target model such that the attack’s success rate is comparable to its success rate when using the Full Grid. For these experiments, we use $T=20$ HPO trials. Figure 9 shows that for matching the performance of the Full Grid ($C=N=M$), the attackers needs to have $C\geq 16$ candidate hyperparameters and $N\geq 1$ models trained using each of the candidates.
@@ 6 Empirical Privacy Leakage Due To HPO: Methods
Fig. 10: The plot above compares average MIA vulnerability ($\overline{\textsc{tpr}}$) when the HPO is done using the training data (TD-HPO) against the vulnerability of models for which the HPO is done using external data set (ED-HPO) for different experiments and privacy levels ($\varepsilon=\infty,8,1$) at fixed fpr. The error bars denote the Clopper–Pearson confidence intervals for the tpr along the x and y-axis. If there is no significant additional leakage observed for TD-HPO, the points are expected to fall on the diagonal representing equal privacy leakage for TD- and ED-HPO. If TD-HPO is a more vulnerable setting, the points are expected to fall on the right of the diagonal. Most of the points align with the diagonal showing no significant difference between MIA vulnerability for the two settings.
Next, we study the impact of HPO under DP on the MIA vulnerability of the final released classification model under the assumption that the hyperparameters are not released.
@@
$\begin{split}\eta^{e}_{i}=\texttt{HPO}(\mathcal{D}^{e}_{i}),\\ \mathcal{M}^{e}_{\mathcal{D}_{i},\eta^{e}_{i}}=\texttt{TRAIN}(\mathcal{A},\mathcal{D}_{i},\eta^{e}_{i}),\\ \text{ where }\mathcal{D}^{e}_{i}\cap\mathcal{D}_{j}=\emptyset\quad\forall i,j.\end{split}$ (11)
@@ 7 Empirical Privacy Leakage Due To HPO: Results
Figure 10 summarizes the average MIA vulnerability for both TD-HPO and ED-HPO for different attacks at varying levels of privacy ($\varepsilon=\infty,8,1$) with $\delta=10^{-5}$. The vulnerability to MIA is very similar between TD and ED at all fpr levels, as observed by most of the points in Figure 10 being on or near the diagonal representing $\overline{\textsc{tpr}}_{\text{TD}}=\overline{\textsc{tpr}}_{\text{ED}}$. For the points farther from the diagonal, the error bars (representing the Clopper–Pearson confidence interval) are large and overlap with the diagonal. In Figure 10, We provide a granular version of these plots for each $\varepsilon$ in Figure A6 and Figure A7.
@@
Figure 11 shows the $p$-values of different tests and the corresponding BY-adjusted p-values on a logarithmic scale sorted in ascending order. We can see that the number of rejections of $H_{0}$ after controlling the FDR drops from three to zero for the paired t-test. This suggests that among our experiments, we do not see any examples of TD-HPO resulting in a statistically significant increase of MIA vulnerability compared to the ED-HPO. We observe exactly the same using permutation test: without FDR-control we reject the null three times and after adjusting we retain the null for all the tests (see Figure A8). Interestingly, no DP experiment yields significant $p$-values even without the adjustment, as all the most significant cases are with non-DP models.